OpenStudy (anonymous):

What is the solution to the equation x – 12 = –3? A. x = 15 B. x = 9 C. x = –15 D. x = –9

OpenStudy (anonymous):

C

OpenStudy (anonymous):

it is..

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yes it is b. 9

OpenStudy (haseeb96):

x -12 =-3 x = -3 +12 x = ??? this is your job to solve this

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

your welcome